# From tug-of-war to Brownian Boost: explicit ODE solutions for player-funded stochastic-differential games

### v3

@@ 4.5.2. Overview of the proof at large
Since $\mathsf{Trap}$, $\mathsf{ProxyTrap}\setminus\mathsf{Trap}$ and $\mathsf{ProxyTrap}^{c}$ partition the space of outcomes, we may add (40) and the two bounds that follow it to obtain
$\mathbb{E}^{i}\Big[P_{-}(S_{-}^{\mathrm{dev}}[\mathrm{alt}],S_{+})\Big]\geq\mathbb{E}^{i}\Big[P_{-}(S_{-}^{\mathrm{dev}},S_{+})\Big]+\big(n_{\infty}-n_{*}\big)\mathbb{P}^{i}(\mathsf{Trap})-\big(n_{\infty,-\infty}\big)\epsilon\,.$
Choosing $\epsilon$ to be less than $\tfrac{n_{\infty}-n_{*}}{n_{\infty,-\infty}}\mathbb{P}^{i}(\mathsf{Trap})$, we find that $\mathbb{E}^{i}\big[P_{-}(S_{-}^{\mathrm{dev}}[\mathrm{alt}],S_{+})\big]>\mathbb{E}^{i}\big[P_{-}(S_{-}^{\mathrm{dev}},S_{+})\big]$, as claimed in (37). This completes the proof of Proposition 4.18. ∎
5. Brownian Boost
One of the main themes of this article is that time-homogeneous Markov-perfect Nash equilibria in $\rho$-Brownian Boost are governed by solutions of the ODE pair in Definition 1.1. Here we study the ODE pair and its solutions. In the first subsection, we offer a heuristic explanation for the appearance of the pair, deriving the equations by a formal argument that is applied directly, in continuous time, to $\mathrm{BB}(\rho)$. In the second subsection, we solve the ODE pair analytically, proving Theorem 1.3 (which characterises the solutions explicitly), Proposition 1.5 (which describes the solutions’ behaviour) and other analytic facts needed in Section 6 to understand low-$\kappa$ $\mathrm{TLP}(\kappa,\rho)$.
@@ 5.1. Coupled HJB equations for Brownian Boost
$\frac{a^{\rho}-b^{\rho}}{a^{\rho}+b^{\rho}}m^{\prime}+\frac{m^{\prime\prime}}{2}-a=0\,.$ (41)
@@
$\frac{a^{\rho}-b^{\rho}}{a^{\rho}+b^{\rho}}n^{\prime}+\frac{n^{\prime\prime}}{2}-b=0\,.$ (42)
This is a pair of Markovian forward equations, valid for any stake profile pair $(a,b)$. As we show next, a further equation pair arises from the consideration that $(a,b)$ is a Nash equilibrium. The stability under unilateral deviation manifest in this concept is gauged in terms of mean total net receipt, with the class of perturbed strategies being broader than time-invariant ones. Indeed, let $z\in[0,\infty)$, and suppose that Maxine stakes at rate $z$ during $0,{\rm d}t]$, after which she reverts to the dictates of the stake profile $a:\mathbb{R}\to[0,\infty)$. Writing $m(x,z)$ for her mean net receipt when she plays against Mina’s stake profile $b$, we have that
@@
$a=\frac{2\rho f^{1+\rho}g^{\rho}}{\big(f^{\rho}+g^{\rho}\big)^{2}}\,\,\,\,\textrm{and}\,\,\,\,b=\frac{2\rho f^{\rho}g^{1+\rho}}{\big(f^{\rho}+g^{\rho}\big)^{2}}\,.$ (45)
@@
Revisiting (41) and (42) with these inferences and notation,
@@
Substituting these stake profile formulas into (45) yields
@@ 5.2. Proving properties of the ODE pair
Here we prove Theorem 1.3 and Proposition 1.5. We also derive further information on $\mathrm{BB}(\rho)$ ODE pair solutions in Proposition 5.4. This includes the key identity $\int_{\mathbb{R}}f_{\rho}(x,u)\,{\rm d}u=\int_{\mathbb{R}}g_{\rho}(x,u)\,{\rm d}u$: that is, $\lambda_{\rm max}(0,\rho)=1$, so that no incentive asymmetry may exist at equilibrium. This information will be central to deriving Theorem 1.22 on low-$\kappa$ $\lambda_{\rm max}(\kappa,\rho)$ in Section 6.
Theorem 1.3 classifies default solutions of the $\mathrm{BB}(\rho)$ ODE pair. These solutions are everywhere positive, but we note in passing that in fact all non-negative solutions of (2) are readily classified by use of this theorem. To see this, first note a scaling: if $(f,g)$ is a non-negative solution of the equation pair, then so is $(af,ag)$, for any $a\geq 0$. And if one or other of $f$ and $g$ vanishes at some point, then the function in question is identically zero, by the Picard-Lindelöf theorem. If $f\equiv 0$ then $g(u)=A\,e^{2u}$ for some $A\in[0,\infty)$, while if $g\equiv 0$ then $f(u)=A\,e^{-2u}$. Thus, by Theorem 1.3, the space of non-negative solutions $(f,g)$ of (2) consists of dilations $(af,ag)$ of default solutions $(f,g)$ by factors $a\in[0,\infty)$, and the solutions $\big(A\,e^{-2u},0\big)$ and $\big(0,A\,e^{2u}\big)$ for $A\in[0,\infty)$.
@@
For $\rho,x\in(0,\infty)$, set $J(u):=S_{\rho}(x,u)^{\rho}$, where $S_{\rho}(x,\cdot)$ is specified in Definition 1.2. Then $J$ is the unique solution of the differential equation
$\frac{{\rm d}J(u)}{{\rm d}u}\,=\,-8\rho^{2}\,\frac{J(u)^{2}}{\big(1+J(u)\big)^{2}}\,\,\,\,\textrm{with}\,\,\,\,J(0)=x^{\rho}\,.$
Remarks: (1). This result has the consequence that $S_{\rho}(x,u)^{\rho}=S_{1}(x^{\rho},\rho^{2}u)$, since the right-hand expression is also a solution of the equation.
(2). Integrating the equation, we find that $J(u)^{2}e^{J(u)-J(u)^{-1}}=e^{-8\rho^{2}u}$ when $x=1$. In view then of what we just noted,
@@
Proof of Lemma 5.1. The initial condition $J(0)=S_{\rho}(x,0)^{\rho}=x^{\rho}$ holds. Differentiating $J(u)=S_{\rho}(x,u)^{\rho}$ gives
$J^{\prime}(u)\,=\,\rho S_{\rho}(x,u)^{\rho-1}S_{\rho}^{\prime}(x,u)\,=\,\rho S_{\rho}^{\rho-1}\times\frac{-8\rho\,S_{\rho}^{1+\rho}}{(1+S_{\rho}^{\rho})^{2}}\,=\,-8\rho^{2}\frac{J(u)^{2}}{(1+J(u))^{2}}\,$,
@@
Let $\rho\in(0,\infty)$. For a pair of differentiable functions $f,g:\mathbb{R}\to(0,\infty)$, set $\phi_{f}=\frac{f^{\prime}}{2f}$, $\phi_{g}=-\frac{g^{\prime}}{2g}$ and $j=(g/f)^{\rho}$. The pair $(f,g)$ is a solution of (2) if and only if the pair of equations
$\big(\phi_{f},\phi_{g}\big)\,=\,\bigg(\frac{2\rho j-(1-j^{2})}{(1+j)^{2}},\frac{2\rho j+(1-j^{2})}{(1+j)^{2}}\bigg)\,.$
@@
Proof. Divide the first equation in the pair (2) by $f$ and write in terms of $F:=f^{\rho}$ and $H:=g^{\rho}$ to obtain
$2\rho FH=F^{2}-H^{2}+\frac{1}{2}\frac{f^{\prime}}{f}(F+H)^{2}=F^{2}-H^{2}+\phi_{f}(F+H)^{2}\,$,
where we use $\phi_{f}=\frac{f^{\prime}}{2f}$. Thus the pair $(f,g)$ satisfies the first equation in (2) if and only if
$\phi_{f}=\frac{2\rho FH-(F^{2}-H^{2})}{(F+H)^{2}}\,$,
@@
$\phi_{f}=\frac{2\rho j-(1-j^{2})}{(1+j)^{2}}\,$, (46)
where we have introduced the function $j=H/F$ after dividing by the positive $F^{2}$.
Now divide the second equation (2) by $g$ to find that $2\rho FH$ equals $H^{2}-F^{2}+\phi_{g}(F+H)^{2}$. Again dividing by $F^{2}>0$, we see that $(f,g)$ satisfies this second equation precisely when $\phi_{g}=\frac{2\rho j+(1-j^{2})}{(1+j)^{2}}$.
@@
Proof of Theorem 1.3. Let $x\in(0,\infty)$. In shorthand, we will denote $f_{\rho}(\cdot)=f_{\rho}(x,\cdot)$, $g_{\rho}(\cdot)=g_{\rho}(x,\cdot)$ and $S(u)=S_{\rho}(x,u)$. Write $F_{\rho}=f_{\rho}(\cdot)^{\rho}$ and $H_{\rho}=g_{\rho}(\cdot)^{\rho}$, and note that these functions are everywhere positive.
We will show that $(f_{\rho},g_{\rho})$ solves (2). To this end, note that, by Definition 1.2,
$f_{\rho}(r)=\exp\!\Bigg\{2\int_{0}^{r}\Phi_{f}(u)\,du\Bigg\},\qquad g_{\rho}(r)=x\cdot\exp\!\Bigg\{-2\int_{0}^{r}\Phi_{g}(u)\,du\Bigg\}\,.$ (47)
@@
$\Phi_{f}=1-\frac{2\big(1+(1-\rho)S^{\rho}\big)}{(1+S^{\rho})^{2}},\qquad\Phi_{g}=1-\frac{2\big((1-\rho)S^{\rho}+S^{2\rho}\big)}{(1+S^{\rho})^{2}}.$ (48)
Moreover, from these expressions for $f_{\rho}$ and $g_{\rho}$, we see that $\Phi_{f}=\frac{f^{\prime}_{\rho}}{2f_{\rho}}$ and $\Phi_{g}=-\frac{g_{\rho}^{\prime}}{2g_{\rho}}$. Differentiating $F_{\rho}=f_{\rho}^{\rho}$ and $H_{\rho}=g_{\rho}^{\rho}$, we also have that
$F_{\rho}^{\prime}=2\rho F_{\rho}\Phi_{f}\,\,\,\,\textrm{and}\,\,\,\,H_{\rho}^{\prime}=-2\rho H_{\rho}\Phi_{g}\,.$ (49)
In order to argue that $(f,g)=(f_{\rho},g_{\rho})$ solves (2), we write $j_{\rho}=H_{\rho}/F_{\rho}$ for the $j$-function attached to the pair $(f_{\rho},g_{\rho})$.
@@
    And also that $j_{\rho}(u)=S_{\rho}(x,u)^{\rho}$ for all $u\in\mathbb{R}$.
Proof: (1,2). Since $g_{\rho}(0)=x$ and $f_{\rho}(0)=1$, $j_{\rho}(0)=x^{\rho}$. By (49),
$\frac{{\rm d}j_{\rho}(u)}{{\rm d}u}\;=\;\frac{H_{\rho}^{\prime}}{F_{\rho}}-\frac{H_{\rho}F_{\rho}^{\prime}}{F_{\rho}^{2}}=-2\rho j_{\rho}\big(\Phi_{f}+\Phi_{g}\big)\,.$ (50)
Writing $J=S^{\rho}>0$ as in Lemma 5.1, we find from (48) that
@@
whence Lemma 5.3(1) holds. Returning to (50), we obtain Lemma 5.3(2).
(3). By Lemma 5.1, $J=S^{\rho}$ from satisfies the differential equation in that result and may be compared to the solution $j_{\rho}$ of the related differential equation in the preceding part. Consequently, $j_{\rho}^{\prime}J=j_{\rho}J^{\prime}$. Consider the ratio $q(u)=j_{\rho}(u)/J(u)$. The derivative is a fraction whose denominator is $J^{2}>0$ and whose numerator vanishes by the just obtained identity. So $q^{\prime}=0$ identically. Thus $q=1$ since $q(0)=j_{\rho}(0)/J(0)=x^{\rho}/x^{\rho}=1$. Hence $j_{\rho}=J=S^{\rho}$ and we obtain Lemma 5.3(3). ∎
By Lemma 5.2, we may prove that $(f_{\rho},g_{\rho})$ solves (2) by showing that
@@
To prove the converse direction in this theorem, let $(f,g)$ be a default solution of (2), so that $f(0)=1$ and $g(0)>0$. By Lemma 5.2, the pair $(\phi_{f},-\phi_{g})$ of one-half logarithmic derivatives satisfies
$\phi_{f}=\frac{2\rho j-(1-j^{2})}{(1+j)^{2}}\,\,\,\,\textrm{and}\,\,\,\,\phi_{g}=\frac{2\rho j+(1-j^{2})}{(1+j)^{2}}\,\,\,\,\textrm{with}\,\,\,\,j=(g/f)^{\rho}\,$, (51)
whence $\phi_{f}+\phi_{g}=4\rho j/(1+j)^{2}$. But $j^{\prime}/j=-2\rho\big(\phi_{f}+\phi_{g}\big)$, so that $j^{\prime}=-8\rho^{2}j^{2}/(1+j)^{2}$. Note that $j(0)=x^{\rho}$ where we set $x=g(0)>0$. Thus $j$ solves the initial value problem satisfied by $J(u)=S_{\rho}(x,u)^{\rho}$ in Lemma 5.1. By the uniqueness claim in this lemma, $j=J$. Hence, $j(u)=J(u)=S_{\rho}(x,u)^{\rho}$ for all $u\in\mathbb{R}$. Since $\phi_{f}$ and $\phi_{g}$ are one-half logarithmic derivatives, we have
$f_{\rho}(x,r)=f(0)\cdot\exp\Big\{2\int_{0}^{r}\phi_{f}(u)\,du\Big\}\,\,\,\,\textrm{and}\,\,\,\,g_{\rho}(x,r)=g(0)\cdot\exp\Big\{-2\int_{0}^{r}\phi_{g}(u)\,du\Big\}\,.$
Now (51) alongside $j=S_{\rho}^{\rho}$ exhibits the pair $(f,g)$ in the desired form $\big(f_{\rho}(x,\cdot),g_{\rho}(x,\cdot)\big)$. The converse direction thus treated, this completes the proof of Theorem 1.3. ∎
@@
    Now let $v=v(x)$ denote the unique real number such that $8\rho v=2\log x+\rho^{-1}\big(x^{\rho}-x^{-\rho}\big)$. Then $v$ is the unique solution of $S_{\rho}(x,v)=1$, and
    $f_{\rho}(1,r)=\frac{f_{\rho}(x,v+r)}{f_{\rho}(x,v)}\,\,\,\,\textrm{and}\,\,\,\,g_{\rho}(1,r)=\frac{g_{\rho}(x,v+r)}{g_{\rho}(x,v)}\,.$

### v4

@@ 4.5.2. Overview of the proof at large
Since $\mathsf{Trap}$, $\mathsf{ProxyTrap}\setminus\mathsf{Trap}$ and $\mathsf{ProxyTrap}^{c}$ partition the space of outcomes, we may add (4.14) and the two bounds that follow it to obtain
$\mathbb{E}^{h}\Big[P_{-}(S_{-}^{\mathrm{dev}}[\mathrm{alt}],S_{+})\Big]\geq\mathbb{E}^{h}\Big[P_{-}(S_{-}^{\mathrm{dev}},S_{+})\Big]+\big(n_{\infty}-n_{*}\big)\mathbb{P}^{h}(\mathsf{Trap})-\big(n_{\infty,-\infty}\big)\epsilon\,.$
Choosing $\epsilon$ to be less than $\tfrac{n_{\infty}-n_{*}}{n_{\infty,-\infty}}\mathbb{P}^{h}(\mathsf{Trap})$, we find that $\mathbb{E}^{h}\big[P_{-}(S_{-}^{\mathrm{dev}}[\mathrm{alt}],S_{+})\big]>\mathbb{E}^{h}\big[P_{-}(S_{-}^{\mathrm{dev}},S_{+})\big]$, as claimed in (4.11). This completes the proof of Proposition 4.18. ∎
Part II Back to Brownian Boost
Chapter 5 The Brownian Boost ODE pair
This part returns to Brownian Boost and develops its analysis in two chapters. In the present chapter, we begin with a heuristic explanation for why time-invariant equilibria of the game are governed by solutions of the $\mathrm{BB}(\rho)$ ODE pair in Definition 1.13, and we then undertake a detailed analytic study of these solutions. In the following chapter, Brownian Boost is realised as a fine-mesh, high-noise limit of the Trail of Lost Pennies, leading to the low-$\kappa$ $\lambda_{\rm max}$ Theorem 1.23 and the asymptotic stakes-and-gameplay Theorem 1.18.
The present chapter is divided into two sections. The first offers a formal but heuristic derivation of the $\mathrm{BB}(\rho)$ ODE pair, obtained by analysing $\mathrm{BB}(\rho)$ directly in continuous time. The second solves this ODE system explicitly, establishing Theorem 1.15, describing the qualitative behaviour of solutions in Proposition 1.19, and recording further analytic properties that will be used in Chapter 6 to analyse the low-$\kappa$ regime of $\mathrm{TLP}(\kappa,\rho)$.
@@ 5.1. Coupled HJB equations for Brownian Boost
$\frac{a^{\rho}-b^{\rho}}{a^{\rho}+b^{\rho}}m^{\prime}+\frac{m^{\prime\prime}}{2}-a\,=\,0\,.$ (5.1)
@@
$\frac{a^{\rho}-b^{\rho}}{a^{\rho}+b^{\rho}}n^{\prime}+\frac{n^{\prime\prime}}{2}-b\,=\,0\,.$ (5.2)
This is a pair of Markovian dynamic programming equations, valid for any stake profile pair $(a,b)$. As we show next, a further equation pair arises from the consideration that $(a,b)$ is a Nash equilibrium. The stability under unilateral deviation manifest in this concept is gauged in terms of mean total net receipt, with the class of perturbed strategies being broader than time-invariant ones. Indeed, let $z\in[0,\infty)$, and suppose that Maxine stakes at rate $z$ during $0,{\rm d}t]$, after which she reverts to the dictates of the stake profile $a:\mathbb{R}\to[0,\infty)$. Writing $m(x,z)$ for her mean net receipt when she plays against Mina’s stake profile $b$, we have that
@@
$a=\frac{2\rho f^{1+\rho}g^{\rho}}{\big(f^{\rho}+g^{\rho}\big)^{2}}\,\,\,\,\textrm{and}\,\,\,\,b=\frac{2\rho f^{\rho}g^{1+\rho}}{\big(f^{\rho}+g^{\rho}\big)^{2}}\,.$ (5.5)
@@
Revisiting (5.1) and (5.2) with these inferences and notation,
@@
Substituting these stake profile formulas into (5.5) yields
@@
Remark. This formal derivation of the $\mathrm{BB}(\rho)$-ODE pair is valid for $\rho\in(0,\infty)$. On the discrete side, the region $W$ of $(\kappa,\rho)$-space specified in (1.7) and utilized in computing ABMN($\kappa,\rho$) elements contains in its boundary the whole ‘Brownian Boost line’ $\{0\}\times(0,\infty)$. These indications may suggest that discrete solution impediments lose their force in the low $\kappa$ limit, and that $\mathrm{BB}(\rho)$ equilibria may be computed for all $\rho\in(0,\infty)$.
These impressions are misleading: the derivation above identifies critical points of the players’ optimization problems for all $\rho\in(0,\infty)$, but it does not guarantee their global optimality. As noted in Section 2.5, the one-step game Penny Forfeit admits a global maximizer in pure strategies throughout the regime $\rho\leq 1$, while for $\rho>1$ global optimality may fail, independently of the value of $\kappa\in(0,1)$. Accordingly, our rigorous results for Brownian Boost are confined to $\rho\in(0,1]$, and the analysis of higher values of $\rho$ remains open.
Moscarini and Smith’s game
In [39] a continuous-time tug-of-war game in which the counter’s drift is given by $a(X_{t})^{1/2}-b(X_{t})^{1/2}$ is studied—in our terms, the game is Brownian Boost with $\rho=1/2$ and the drift denominator omitted, so that the drift term in (1) becomes homogeneous in the stake pair $(a,b)$ with degree one-half rather than zero. Moscarini and Smith focus on Markov-perfect equilibria for games played on finite intervals centred at the origin, restricting attention to symmetric equilibria, in which reflection through the origin corresponds to interchange of the players’ continuation values; they obtain explicit characterisations of such equilibria.
Their analysis may be recast in a form closely parallel to the heuristic HJB derivation given above for Brownian Boost. Introducing value derivatives $f=m^{\prime}>0$ and $g=-n^{\prime}>0$, the equilibrium conditions lead to the coupled ODE system
$f^{\prime}=f\big(g-f/2\big),\qquad g^{\prime}=g\big(g/2-f\big),$
with accompanying stake rates $a=f^{2}/4$ and $b=g^{2}/4$. By translation invariance, one may take $f(0)=g(0)>0$.
Analysing the MS-ODE pair shows that for positive arguments $g<f$ and $g/f\to 0$, so a discouragement effect is present. Moreover, $f^{\prime}(u)\sim-f(u)^{2}/2$ and $f(u)\sim 2/u$ for large $u$. As a consequence, $\int_{\mathbb{R}}f(r)\,{\rm d}r=\int_{\mathbb{R}}g(r)\,{\rm d}r=\infty$, and thus no MS equilibria on the real line exist with finite terminal prizes. This necessitates the restriction, adopted in [39], to finite gameboards.
Although Moscarini and Smith recover a discouragement phenomenon akin to that identified by Harris and Vickers, there does not appear to be a counterpart of the delicate $\lambda_{\rm max}=1$ and $\lambda_{\rm max}-1\in[0,10^{-4}]$ effects that we find rigorously and numerically in $\mathrm{BB}(\rho)$ and $\mathrm{TLP}(\kappa,\rho)$. Indeed, $\lambda_{\rm max}$ may naturally be interpreted as the supremum, over finite gameboards $I$ and solutions $(f,g)$ of the above ODE pair, of the statistic $\int_{I}g(r)\,{\rm d}r/\int_{I}f(r)\,{\rm d}r$. Gameboards must be finite intervals in order for these quantities to be well-defined positive real values, but variation over all such intervals simply yields $\lambda_{\rm max}=\infty$, since $g\gg f$ for $u\ll 0$.
Nor does a sharply localised battlefield region emerge: on long intervals, equilibrium stakes and continuation values vary relatively slowly across the interior. It would be interesting to extend Moscarini and Smith’s classification beyond the symmetric case. From the ODE perspective, one natural approach is to fix a solution $(f,g)$ with $f(0)=g(0)$ and associate to any finite interval $I\subset\mathbb{R}$ the triple $\big(\int_{I}f,\,\int_{I}g,\,|I|\big)$. Determining whether this correspondence is bijective is an open question whose resolution would yield a concrete parametrisation of finite-interval MS equilibria in terms of terminal payoffs and gameboard size.
@@ 5.2. Proving properties of the ODE pair
Here we prove Theorem 1.15 and Proposition 1.19. We also derive further information on $\mathrm{BB}(\rho)$ ODE pair solutions in Proposition 5.4. This includes the key identity $\int_{\mathbb{R}}f_{\rho}(\xi,u)\,{\rm d}u=\int_{\mathbb{R}}g_{\rho}(\xi,u)\,{\rm d}u$, valid for any choice of the flow index $\xi\in(0,\infty)$: in the notation of  (1.18), $\lambda_{\rm max}(0,\rho)=1$, so that no incentive asymmetry may exist at equilibrium. This information will be central to deriving Theorem 1.23 on low-$\kappa$ $\lambda_{\rm max}(\kappa,\rho)$ in Chapter 6.
Theorem 1.15 classifies default solutions of the $\mathrm{BB}(\rho)$ ODE pair and, as we now note in passing, it thereby also classifies all strictly positive solutions. Indeed, the dilation $(f,g)\mapsto(af,ag)$ preserves the positive solution space for any $a>0$, so that any such solution is equivalent, up to dilation, to a default one. It follows immediately from Theorem 1.15 that the strictly positive solutions are precisely the positive dilations of default solutions. Note also that $\big(A\,e^{-2u},0\big)$ and $\big(0,A\,e^{2u}\big)$, for $A\in[0,\infty)$, are non-negative solutions.
@@
For $\rho,\xi\in(0,\infty)$, set $J(u):=S_{\rho}(\xi,u)^{\rho}$, where $S_{\rho}(\xi,\bullet)$ is specified in Definition 1.14. Then $J$ is the unique solution of the differential equation
$\frac{{\rm d}J(u)}{{\rm d}u}\,=\,-8\rho^{2}\,\frac{J(u)^{2}}{\big(1+J(u)\big)^{2}}\,\,\,\,\textrm{with}\,\,\,\,J(0)=\xi^{\rho}\,.$
Remarks: (1). This result has the consequence that $S_{\rho}(\xi,u)^{\rho}=S_{1}(\xi^{\rho},\rho^{2}u)$, since the right-hand expression is also a solution of the equation.
(2). Integrating the equation, we find that $J(u)^{2}e^{J(u)-J(u)^{-1}}=e^{-8\rho^{2}u}$ when $\xi=1$. In view then of what we just noted,
@@
Proof of Lemma 5.1. The initial condition $J(0)=S_{\rho}(\xi,0)^{\rho}=\xi^{\rho}$ holds. Differentiating $J(u)=S_{\rho}(\xi,u)^{\rho}$ gives
$J^{\prime}(u)\,=\,\rho S_{\rho}(\xi,u)^{\rho-1}S_{\rho}^{\prime}(\xi,u)\,=\,\rho S_{\rho}^{\rho-1}\times\frac{-8\rho\,S_{\rho}^{1+\rho}}{(1+S_{\rho}^{\rho})^{2}}\,=\,-8\rho^{2}\frac{J(u)^{2}}{(1+J(u))^{2}}\,$,
@@
Let $\rho\in(0,\infty)$. For a pair of differentiable functions $f,g:\mathbb{R}\to(0,\infty)$, set $\phi_{f}=\frac{f^{\prime}}{2f}$, $\phi_{g}=-\frac{g^{\prime}}{2g}$ and $j=(g/f)^{\rho}$. The pair $(f,g)$ is a solution of (1.13) if and only if the pair of equations
$\big(\phi_{f},\phi_{g}\big)\,=\,\bigg(\frac{2\rho j-(1-j^{2})}{(1+j)^{2}},\frac{2\rho j+(1-j^{2})}{(1+j)^{2}}\bigg)$
@@
Proof. Divide the first equation in the pair (1.13) by $f$ and write in terms of $F:=f^{\rho}$ and $G:=g^{\rho}$ to obtain
$2\rho FG=F^{2}-G^{2}+\frac{1}{2}\frac{f^{\prime}}{f}(F+G)^{2}=F^{2}-G^{2}+\phi_{f}(F+G)^{2}\,$,
where we use $\phi_{f}=\frac{f^{\prime}}{2f}$. Thus the pair $(f,g)$ satisfies the first equation in (1.13) if and only if
$\phi_{f}=\frac{2\rho FG-(F^{2}-G^{2})}{(F+G)^{2}}\,$,
@@
$\phi_{f}=\frac{2\rho j-(1-j^{2})}{(1+j)^{2}}\,$, (5.6)
where we have introduced the function $j=G/F$ after dividing by the positive $F^{2}$.
Now divide the second equation (1.13) by $g$ to find that $2\rho FG$ equals $G^{2}-F^{2}+\phi_{g}(F+G)^{2}$. Again dividing by $F^{2}>0$, we see that $(f,g)$ satisfies this second equation precisely when $\phi_{g}=\frac{2\rho j+(1-j^{2})}{(1+j)^{2}}$.
@@
Proof of Theorem 1.15. Let $\xi\in(0,\infty)$. In shorthand, we will denote $f_{\rho}(\bullet)=f_{\rho}(\xi,\bullet)$, $g_{\rho}(\bullet)=g_{\rho}(\xi,\bullet)$ and $S(\bullet)=S_{\rho}(\xi,\bullet)$. Write $F_{\rho}=f_{\rho}(\bullet)^{\rho}$ and $G_{\rho}=g_{\rho}(\bullet)^{\rho}$, and note that these functions are everywhere positive.
We will show that $(f_{\rho},g_{\rho})$ solves (1.13). To this end, note that, by Definition 1.14,
$f_{\rho}(u)=\exp\!\Bigg\{2\int_{0}^{u}\Phi_{f}(r)\,{\rm d}r\Bigg\},\qquad g_{\rho}(u)=\xi\cdot\exp\!\Bigg\{-2\int_{0}^{u}\Phi_{g}(r)\,{\rm d}r\Bigg\}\,.$ (5.7)
@@
$\Phi_{f}=1-\frac{2\big(1+(1-\rho)S^{\rho}\big)}{(1+S^{\rho})^{2}},\qquad\Phi_{g}=1-\frac{2\big((1-\rho)S^{\rho}+S^{2\rho}\big)}{(1+S^{\rho})^{2}}.$ (5.8)
Moreover, from these expressions for $f_{\rho}$ and $g_{\rho}$, we see that $\Phi_{f}=\frac{f^{\prime}_{\rho}}{2f_{\rho}}$ and $\Phi_{g}=-\frac{g_{\rho}^{\prime}}{2g_{\rho}}$. Differentiating $F_{\rho}=f_{\rho}^{\rho}$ and $G_{\rho}=g_{\rho}^{\rho}$, we also have that
$F_{\rho}^{\prime}=2\rho\,F_{\rho}\Phi_{f}\,\,\,\,\textrm{and}\,\,\,\,G_{\rho}^{\prime}=-2\rho\,G_{\rho}\Phi_{g}\,.$ (5.9)
In order to argue that $(f,g)=(f_{\rho},g_{\rho})$ solves (1.13), we write $j_{\rho}=G_{\rho}/F_{\rho}$ for the $j$-function attached to the pair $(f_{\rho},g_{\rho})$.
@@
    And also that $j_{\rho}(u)=S_{\rho}(\xi,u)^{\rho}$ for all $u\in\mathbb{R}$.
Proof: (1,2). Since $g_{\rho}(0)=\xi$ and $f_{\rho}(0)=1$, $j_{\rho}(0)=\xi^{\rho}$. By (5.9),
$\frac{{\rm d}j_{\rho}(u)}{{\rm d}u}\;=\;\frac{G_{\rho}^{\prime}}{F_{\rho}}-\frac{G_{\rho}F_{\rho}^{\prime}}{F_{\rho}^{2}}=-2\rho j_{\rho}\big(\Phi_{f}+\Phi_{g}\big)\,.$ (5.10)
Writing $J=S^{\rho}>0$ as in Lemma 5.1, we find from (5.8) that
@@
whence Lemma 5.3(1) holds. Returning to (5.10), we obtain Lemma 5.3(2).
(3). By Lemma 5.1, $J=S^{\rho}$ from satisfies the differential equation in that result and may be compared to the solution $j_{\rho}$ of the related differential equation in the preceding part. Consequently, $j_{\rho}^{\prime}J=j_{\rho}J^{\prime}$. Consider the ratio $q(u)=j_{\rho}(u)/J(u)$. The derivative is a fraction whose denominator is $J^{2}>0$ and whose numerator vanishes by the just obtained identity. So $q^{\prime}=0$ identically. Thus $q=1$ since $q(0)=j_{\rho}(0)/J(0)=\xi^{\rho}/\xi^{\rho}=1$. Hence $j_{\rho}=J=S^{\rho}$ and we obtain Lemma 5.3(3). ∎
By Lemma 5.2, we may prove that $(f_{\rho},g_{\rho})$ solves (1.13) by showing that
@@
To prove the converse direction in this theorem, let $(f,g)$ be a default solution of (1.13), so that $f(0)=1$ and $g(0)>0$. By Lemma 5.2, the pair $(\phi_{f},-\phi_{g})$ of one-half logarithmic derivatives satisfies
$\phi_{f}=\frac{2\rho j-(1-j^{2})}{(1+j)^{2}}\,\,\,\,\textrm{and}\,\,\,\,\phi_{g}=\frac{2\rho j+(1-j^{2})}{(1+j)^{2}}\,\,\,\,\textrm{with}\,\,\,\,j=(g/f)^{\rho}\,$, (5.11)
whence $\phi_{f}+\phi_{g}=4\rho j/(1+j)^{2}$. But $j^{\prime}/j=-2\rho\big(\phi_{f}+\phi_{g}\big)$, so that $j^{\prime}=-8\rho^{2}j^{2}/(1+j)^{2}$. Note that $j(0)=\xi^{\rho}$ where we set $\xi=g(0)>0$. Thus $j$ solves the initial value problem satisfied by $J(u)=S_{\rho}(\xi,u)^{\rho}$ in Lemma 5.1. By the uniqueness claim in this lemma, $j=J$. Hence, $j(u)=J(u)=S_{\rho}(\xi,u)^{\rho}$ for all $u\in\mathbb{R}$. Since $\phi_{f}$ and $\phi_{g}$ are one-half logarithmic derivatives, we have
$f(u)=f(0)\cdot\exp\Big\{2\int_{0}^{u}\phi_{f}(r)\,{\rm d}r\Big\}\,\,\,\,\textrm{and}\,\,\,\,g(u)=g(0)\cdot\exp\Big\{-2\int_{0}^{u}\phi_{g}(r)\,{\rm d}r\Big\}$
with $\big(f(0),g(0)\big)=(1,\xi)$. Now (5.11) alongside $j=S_{\rho}^{\rho}$ exhibits the pair $(f,g)$ in the desired form $\big(f_{\rho}(\xi,\bullet),g_{\rho}(\xi,\bullet)\big)$. The converse direction thus treated, this completes the proof of Theorem 1.15. ∎
@@
    Now let $v=v(\xi)$ denote the unique real number such that $8\rho v=2\log\xi+\rho^{-1}\big(\xi^{\rho}-\xi^{-\rho}\big)$. The map $(0,\infty)\to\mathbb{R}:\xi\to v$ is an increasing bijection. The value $v=v(\xi)$ is the unique solution of $S_{\rho}(\xi,v)=1$. We have
    $f_{\rho}(1,u)=\frac{f_{\rho}(\xi,v+u)}{f_{\rho}(\xi,v)}\,\,\,\,\textrm{and}\,\,\,\,g_{\rho}(1,u)=\frac{g_{\rho}(\xi,v+u)}{g_{\rho}(\xi,v)}\,.$